Hey i'm having some trouble in math so I would like some help from you guys (:
Which of the following equations represents a line perpendicular to the line y-2x=5

options are
A. x - 1/2y =-10
B. x + 1/2y=10
C. X+2y=10
D -x+2y=10

Answer by mananth(16946)

y-2x=5
..
y = 2x +5
..
the line which has slope of -1/2 is the line perpendicular to the above line
..
option C
x+2y=10
2y=-x+10
/2
y=-x/2+5
slope = -1/2
...
m.ananth@hotmail.ca